OpenStudy (anonymous):

How do i solve this? : -2(x+2)(x-2) ? Thanks.

OpenStudy (anonymous):

by "solve" do you mean "multiply"?

OpenStudy (anonymous):

Yes, this is part of a bigger problem, this part is what throws me off :)

OpenStudy (anonymous):

-2 (x2-4) \[-2x ^{2}\]+8

OpenStudy (anonymous):

-2x2+8

OpenStudy (anonymous):

\[(x+2)(x-2)=x^2-4\] so \[2(x+2)(x-2)=-2x^2+8\]

OpenStudy (anonymous):

\[-2x ^{2}+8\]

OpenStudy (anonymous):

Thank you so much!! :D

OpenStudy (anonymous):

*hugs everyone*